find the work done by the system in reversible process for which PV^n=constant?

Grade:11

1 Answers

Vikas TU
7 years ago
Work done is the integral form for dw in reversible process that is:
dW = Pext. dV
Now we have given as: => PV^n = k
or
P =  kV^-n
Put Pressure into dw eqn. we get,
dW = k(V^-n)dV
Integerating both sides we get,
W = k(V^(-n+1)/(1 – n))
again put value of k this time we get,
W = PV^n(V^(-n+1)/(1 – n)) = > PV/(1 – x) => nRT/(1 – n)